OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 9 and 36. What is the length of the altitude? Answer 3 6 18 45

OpenStudy (anonymous):

@ParthKohli , @AccessDenied please try and help

OpenStudy (anonymous):

18 do you want to know why?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

 now you can see the right triangle, and the numbers are the length of the lines, suppose the altitude's length is 'x',then 9/x=x/36 \[x ^{2}=36\times9\] x=18